need surface area and volume of cylinder with radius of 20 inches and height of 25 inches
Answer by mananth(16946)

need surface area and volume of cylinder with radius of 20 inches and height of 25 inches
volume of cylinder = pi*r^2*h
=22/7 *20*20*25
=31400 cu. inches
surface area of cylinder = 2*pi*r*h
=2*3.14*20*25
3140 square inches